# SVIP: Towards Verifiable Inference of Open-source Large Language Models

### v1

@@ Acknowledgments
Ethics Statement
@@
Reproducibility Statement
Our code repository is available at https://github.com/ASTRAL-Group/SVIP_LLM_Inference_Verification. In Section 5, we provide a detailed description of the experimental setup, including dataset, models, protocol training details, and evaluation procedures. Additional experimental details can be found in Appendix D.
Appendix A Discussions
A.1 Limitations and Future Work
In our SVIP protocol, although the labeling network $y_{\gamma}(\cdot)$ can be applied to multiple specified models once trained, the proxy task head $f_{\phi}(\cdot)$, proxy task feature extractor $g_{\theta}(\cdot)$, and secret embedding network $t_{\psi}(\cdot)$ need to be optimized for each specified model. Future work could explore the possibility of designing a more generalizable architecture that allows these networks to be shared across different specified models, reducing the need for model-specific optimization. Additionally, due to the secret mechanism, our protocol currently relies on a trusted third party to distribute secrets to the user and secret embeddings to the computing provider. Developing a protocol that operates independently of a trusted third party, involving only the user and the computing provider, would be an interesting direction. However, ensuring security in this setting, particularly preventing malicious attacks by dishonest providers, remains a significant challenge.
A.2 Simple Approaches to Verifiable LLM Inference Can Be Vulnerable
One straightforward solution to verifiable LLM inference, as briefly mentioned in Section 1, involves the user curating a small set of prompt examples from established benchmarks and sending them to the computing provider. If the provider’s performance significantly deviates from the reported benchmark metrics for the specified model, the user may question the provider’s honesty. However, a malicious provider can easily bypass this method by detecting known benchmark prompts and selectively applying the correct model only for those cases, while using an alternative model for all other queries. Additionally, testing such benchmark prompts also increases the user’s inference costs.
Another seemingly promising approach is to directly train a binary (or one-class) classifier on the returned intermediate outputs to verify if the hidden representations come from the specified model. However, a simple attack involves the provider caching hidden representations from the correct model that are unrelated to the user’s input. The dishonest provider could then use a smaller LLM for inference and return these cached irrelevant representations to deceive the classifier while saving costs.
A.3 Hypothesis Testing for Verification Using a Batch of Prompt Queries
@@
The null hypothesis assumes that the computing provider is acting honestly, i.e., the specified model is used. Under the null hypothesis, $p$ corresponds to the True Positive Rate (TPR) of our protocol:
$p=\mathbb{P}(V_{i}=1|\mathcal{M}_{spec}~{}\text{is used for inference})=\text{TPR}$ (11)
We define the following test statistic:
$z=\frac{\sum_{i=1}^{B}V_{i}-B\cdot\text{TPR}}{\sqrt{B\cdot\text{TPR}\cdot(1-\text{TPR})}}.$ (12)
For sufficiently large batch sizes ($B\geq 30$, as is common in practice), we approximate the distribution of $z$ as normal. We reject the null hypothesis at a $95\%$ confidence level when $z<-1.645$, indicating strong evidence that the cloud provider is acting dishonestly. In practice, since the true TPR is unknown, we can approximate the TPR using the estimated empirical FNR from Section 5.2, i.e., $\hat{\text{TPR}}=1-\hat{\text{FNR}}$, and the test statistic becomes:
$\hat{z}=\frac{\sum_{i=1}^{B}V_{i}-B\cdot\hat{\text{TPR}}}{\sqrt{B\cdot\hat{\text{TPR}}\cdot(1-\hat{\text{TPR}})}}.$ (13)
Appendix B Extended related work
@@
Open-source LLMs are freely available models that offer flexibility for use and modification. Popular examples include GPT-Neo (Black et al., 2022), BLOOM (Le Scao et al., 2023), Llama (Touvron et al., 2023a; b; Dubey et al., 2024), Mistral (Jiang et al., 2023), and Falcon (Almazrouei et al., 2023). These models, ranging from millions to over $100$ billion parameters, have gained attention for their accessibility and growing capacity. However, larger models like Falcon-40B (Almazrouei et al., 2023), and Llama-3.1-70B (Dubey et al., 2024) come with steep computational costs, making even inference impractical on local machines due to the significant GPU memory required. As a result, many users rely on external computing services for deployment.
Additional Background on Cryptographic VC Techniques
Among cryptographic VC techniques, proof-based methods involve the generation of mathematical proofs that certify the correctness of outsourced computations. Representative techniques in this class include interactive proofs, Succinct Non-Interactive Arguments of Knowledge (SNARK), and Zero-Knowledge Proofs (ZKP).
Interactive proofs involve multiple rounds of interaction between a verifier (the user) and a prover (the computing provider) to ensure the computation’s integrity (Cormode et al., 2011; Goldwasser et al., 2015; Thaler, 2013). SNARK allows a verifier to validate a computation with a single, short proof that requires minimal computational effort (Fiore et al., 2020; Bontekoe et al., 2023). ZKP further enhances privacy by enabling the prover to convince the verifier of a statement’s truth without revealing any additional information beyond the validity of the claim (Fiege et al., 1987; De Santis & Persiano, 1992). Due to their rigorous guarantees of correctness and privacy, these techniques have been widely applied in blockchain and related areas (Yang & Li, 2020; Sun et al., 2021; Šimunić et al., 2021).
Appendix C Additional Attacks
@@
Appendix D Additional Experimental Details and Results
D.1 Datasets and Models
A selection of prompt inputs from the LMSYS-Chat-1M, ToxicChat, and web_questions datasets is presented in Tables 5, 6, and 7, respectively. A moderate distribution shift is observed across these datasets. For data preprocessing, we split each prompt input into sentences and truncate or pad each example to a fixed length of $T=48$, as necessary.
@@ D.1 Datasets and Models
Table 5: Example prompt inputs from the LMSYS-Chat-1M dataset
@@
Table 6: Example prompt inputs from the ToxicChat dataset
@@
Table 7: Example prompt inputs from the web_questions dataset
@@
We select $5$ widely-used LLMs as the specified models in our experiment, including Llama-2-13B (Touvron et al., 2023b), GPT-NeoX-20B (Black et al., 2022), OPT-30B (Zhang et al., 2023), Falcon-40B (Almazrouei et al., 2023), and Llama-3.1-70B (Dubey et al., 2024). As alternative models, we use $6$ smaller LLMs, including GPT2-XL ($1.5$B) (Radford et al., 2019), GPT-NEO-2.7B (Gao et al., 2020), GPT-J-6B (Wang & Komatsuzaki, 2021), OPT-6.7B (Zhang et al., 2022), Vicuna-7B (Zheng et al., 2023b) and Llama-2-7B (Touvron et al., 2023b). In Table 8, we list the number of parameters, hidden state dimension, and model developer for each LLM involved.
@@
D.2 Additional Protocol Training Details
@@
where the first item is the contrastive loss introduced in Eq. (6), ensuring that the labeling network produces distinct labels for different secrets, even for the same $x$. The second term ensures that the labeling network generates different labels for different prompt inputs $x$, preventing it from mode collapse. Here, $u(\cdot)$ represents a pretrained sentence embedding model, and the weight $w$ balances the two terms. We use all-mpnet-base-v2 (Reimers, 2019) as the sentence embedding model and a 2-layer MLP to embed the secret. Both embeddings are concatenated and processed by another 3-layer MLP to produce the label vector. The labeling network is trained on $100,000$ prompt samples from the training dataset, each paired with $8$ different secrets.
@@
D.3 Experimental Details and Additional Results of the Protocol Accuracy
@@
We evaluate the accuracy of our protocol on test web_questions dataset to further assess its generalizability. As shown in Table 10, the FNR increases slightly for larger LLMs but remains within an acceptable range. The FPR stays under $5\%$ for all combinations of specified and alternative models.
Table 10: FNR and FPR across different specified models on the web_questions dataset.
@@
D.4 Examining the Labeling Network
@@
D.5 Experimental Details of Adapter Attack

### v2

@@ Acknowledgments
Appendix A Accessibility
Our code repository is available at https://github.com/ASTRAL-Group/SVIP_LLM_Inference_Verification. In Section 4, we provide a detailed description of the experimental setup, including dataset, models, protocol training details, and evaluation procedures. Additional experimental details can be found in Appendix F.
Appendix B Ethics Statement
@@
Appendix C Discussions
C.1 Limitations and Future Work
In our SVIP protocol, although the labeling network $y_{\gamma}(\cdot)$ can be applied to multiple specified models once trained, the proxy task head $f_{\phi}(\cdot)$, proxy task feature extractor $g_{\theta}(\cdot)$, and secret embedding network $t_{\psi}(\cdot)$ need to be optimized for each specified model. Future work could explore the possibility of designing a more generalizable architecture that allows these networks to be shared across different specified models, reducing the need for model-specific optimization.
Additionally, due to the secret mechanism, our protocol currently relies on the platform to distribute secrets to the user and secret embeddings to the computing provider. Developing a protocol that operates independently of a third party, involving only the user and the computing provider, would be an interesting direction. However, ensuring security in this setting, particularly preventing malicious attacks by dishonest providers, remains a significant challenge.
Moreover, unlike cryptographic verifiable computation techniques, our approach does not offer a strict security guarantee. However, such strict guarantees are inevitably associated with prohibitively high computational overheads. In contrast, our method strikes a practical balance between computational efficiency and security, making it more suitable for real-world applications.
C.2 Hypothesis Testing for Verification Using a Batch of Prompt Queries
A single prompt query may occasionally yield an incorrect verification result due to FNR or FPR. In practice, users often have multiple prompt queries $\{x_{i}\}_{i=1}^{B}$, where $B$ denotes the number of prompts. For each prompt, we observe $V_{i}:=V(x_{i},z(x_{i});\phi^{*},\theta^{*},\psi^{*})\in\{0,1\},i\in[B]$ from Eq. (8).
@@ C.2 Hypothesis Testing for Verification Using a Batch of Prompt Queries
$p_{1}=\mathbb{P}(V_{i}=1\mid\mathcal{M}_{\text{spec}}~{}\text{isused for inference})=\text{TPR}.$ (11)
[…]
In practice, we determine whether the provider is acting honestly based on the mean of the observed values $\{V_{i}\}_{i=1}^{B}$, denoted as
$\bar{V}=\frac{1}{B}\sum_{i=1}^{B}V_{i}.$
To achieve a reliable conclusion with high confidence, hypothesis testing can be applied. Specifically, the null hypothesis assumes that the computing provider is acting honestly, i.e., $Z=1$, and the rejection region is $\bar{V}<\tau$. For sufficiently large numbers of prompt queries ($B\geq 30$, as is common in practice), we adopt a normal approximation to derive the type-I error rate and type-II error rate:
  • Type-I Error Rate ($\alpha$): This is the probability of falsely concluding dishonesty when the provider is honest. Under the null hypothesis ($Z=1$), $\bar{V}\sim\mathcal{N}(p_{1},\frac{p_{1}(1-p_{1})}{B})$. Thus:
    $\alpha=\Phi\left(\frac{\tau-p_{1}}{\sqrt{\frac{p_{1}(1-p_{1})}{B}}}\right),$
    where $\Phi$ denotes the CDF of the standard normal distribution.
  • Type-II Error Rate ($\beta$): This is the probability of falsely concluding honesty when the provider is dishonest. Under the alternative hypothesis ($Z=0$), $\bar{V}\sim\mathcal{N}(p_{0},\frac{p_{0}(1-p_{0})}{B})$. Thus:
    $\beta=1-\Phi\left(\frac{\tau-p_{0}}{\sqrt{\frac{p_{0}(1-p_{0})}{B}}}\right).$
For example, when $p_{0}=0.81\%$ and $p_{1}=1-3.13\%=96.87\%$, corresponding to the case of using Llama-3.1-70B as the specified model and Llama-2-7B as the alternative model (as shown in Table 4.1), with $B=30$, we plot the type-I and type-II error rates under varying thresholds in the range $[0.1,0.9]$.
Figure 6 illustrates that for most thresholds in this range, both the type-I and type-II error rates are significantly smaller than $0.01$, a commonly used strict threshold, and approach zero. For instance, when the threshold is $\tau=0.5$, the type-I and type-II error rates are $1.7\times 10^{-49}$ and $0.0$, respectively. This result demonstrates the strong robustness of our protocol. Further, Figure 7 shows that even with as few as $B=10$ prompt queries, both type-I and type-II error rates remain close to 00 for most thresholds, highlighting the protocol’s reliability with limited samples.
Figure 6: Type-I and type-II error rates under different thresholds. Error rates are below $0.01$ for most thresholds and approach zero.
Figure 7: Type-I and type-II error rates for varying sample sizes ($B=10,20,30$) under different thresholds. Even with $B=10$, both error rates remain below $0.01$ for most thresholds.
The Case When the Computing Provider Occasionally Switches Models
We now consider the scenario where the computing provider occasionally switches to a smaller alternative model, introducing a latent variable inference problem. Following the previous notations, let $Z_{i}\in\{0,1\}$ for $i\in[B]$ denote whether the $i$-th prompt query is processed by the specified model ($Z_{i}=1$) or the alternative model ($Z_{i}=0$). The objective is to infer the unobservable latent states $\{Z_{i}\}_{i=1}^{B}$ based on the observed values $\{V_{i}\}_{i=1}^{B}$. We assume the probability of switching to the smaller model is fixed at $\pi$.
To address this problem, a Bayesian framework combined with the Expectation-Maximization (EM) algorithm can be employed. Using Bayes’ rule, the posterior probability can be expressed as:
$\gamma_{i}:=\mathbb{P}(Z_{i}=1\mid V_{i},p_{1},p_{0},\pi)=\frac{\pi\cdot\mathbb{P}(V_{i}\mid Z_{i}=1;p_{1})}{\pi\cdot\mathbb{P}(V_{i}\mid Z_{i}=1;p_{1})+(1-\pi)\cdot\mathbb{P}(V_{i}\mid Z_{i}=0;p_{0})}.$
Expanding the likelihood terms:
$\gamma_{i}=\frac{\pi\cdot p_{1}^{V_{i}}\cdot(1-p_{1})^{1-V_{i}}}{\pi\cdot p_{1}^{V_{i}}\cdot(1-p_{1})^{1-V_{i}}+(1-\pi)\cdot p_{0}^{V_{i}}\cdot(1-p_{0})^{1-V_{i}}}.$
The parameter updates are derived as:
$p_{1}=\frac{\sum_{i=1}^{B}\gamma_{i}\cdot V_{i}}{\sum_{i=1}^{B}\gamma_{i}},\quad p_{0}=\frac{\sum_{i=1}^{B}(1-\gamma_{i})\cdot V_{i}}{\sum_{i=1}^{B}(1-\gamma_{i})},\quad\pi=\frac{\sum_{i=1}^{B}\gamma_{i}}{B}.$
The EM algorithm iterates between the E-step and M-step until convergence. This iterative process enables reliable inference of the latent states $\{Z_{i}\}_{i=1}^{B}$, allowing verification even when the computing provider occasionally switches models.
C.3 Preservation of Completion Quality
Our protocol requires the computing provider to generate the LLM completion as usual and then additionally return a processed hidden representation for verification. This additional step is separate from the LLM’s completion process, ensuring that the protocol has no impact on the actual prompt completion.
Appendix D Extended Related Work
@@
Open-source LLMs are freely available models that offer flexibility for use and modification. Popular examples include GPT-Neo [2], BLOOM [26], Llama [43, 44, 9], Mistral [19], and Falcon [1]. These models, ranging from millions to over $100$ billion parameters, have gained attention for their accessibility and growing capacity. However, larger models like Falcon-40B [1], and Llama-3.1-70B [9] come with steep computational costs, making even inference impractical on local machines due to the significant GPU memory required. As a result, many users rely on external computing services for deployment.
Verifiable Computing
Verifiable Computing (VC) allows users to verify that an untrusted computing provider has executed computations correctly, without having to perform the computation themselves [46, 50, 7, 22]. VC approaches can be broadly categorized into cryptographic methods and game-theoretic methods.
Cryptographic VC techniques either require the provider to return a mathematical proof that confirms the correctness of the results [14, 38, 34], or rely on secure computation techniques [13, 31, 25]. These techniques cryptographically guarantee correctness and have been applied to machine learning models and shallow neural networks [33, 54, 17, 28, 14, 27]. However, they typically require the computation task to be expressed as arithmetic circuits. Representing open-source LLMs in circuit form is particularly challenging due to their complex architectures and intricate operations. Moreover, the sheer size of these models, with billions of parameters, introduces substantial computational overhead. A recent work, zkLLM [40], attempts to verify LLM inference using Zero Knowledge Proofs. For the Llama-2-13B [44] model, generating a proof for a single prompt takes $803$ seconds, and repeating this process for large batches of prompt queries becomes computationally prohibitive.
Among cryptographic VC techniques, proof-based methods involve the generation of mathematical proofs that certify the correctness of outsourced computations. Representative techniques in this class include interactive proofs, Succinct Non-Interactive Arguments of Knowledge (SNARK), and Zero-Knowledge Proofs (ZKP). Interactive proofs involve multiple rounds of interaction between a verifier (the user) and a prover (the computing provider) to ensure the computation’s integrity [6, 15, 42]. SNARK allows a verifier to validate a computation with a single, short proof that requires minimal computational effort [11, 3]. ZKP further enhances privacy by enabling the prover to convince the verifier of a statement’s truth without revealing any additional information beyond the validity of the claim [10, 8]. Due to their rigorous guarantees of correctness and privacy, these techniques have been widely applied in blockchain and related areas [49, 41, 39].
In contrast, game-theoretic VC techniques ensure the correctness of outsourced computations by leveraging economic incentives to enforce honest behavior [32, 30]. For instance, a sampling-based verification mechanism Proof of Sampling [53] requires multiple computing service providers to independently compute and compare results, ensuring integrity through penalties and rewards. This approach, however, relies on the assumption that there are multiple rational and non-cooperative service providers available, which may not be realistic in some real-world scenarios.
LLM Watermarking and Fingerprinting
LLM watermarking involves embedding algorithmically detectable signals into the text generated by LLMs, with the goal of identifying AI-generated texts [21, 18, 4, 16]. Meanwhile, LLM fingerprinting implants specific backdoor triggers into LLMs, causing the model to generate particular text whenever a confidential private key is used [48]. Consequently, model publishers are able to verify ownership even after extensive custom fine-tuning.
Appendix E Additional Attacks
@@
Appendix F Additional Experimental Details and Results
F.1 Datasets and Models
A selection of prompt inputs from the LMSYS-Chat-1M, ToxicChat, and web_questions datasets is presented in Tables 4, 5, and 6, respectively. A moderate distribution shift is observed across these datasets. For data preprocessing, we filter the datasets to keep only English conversations and extract the user prompts for each conversation. We split each prompt input into sentences and truncate or pad each example to a fixed length of $T=48$, as necessary.
@@ F.1 Datasets and Models
Table 4: Example prompt inputs from the LMSYS-Chat-1M dataset.
@@
Table 5: Example prompt inputs from the ToxicChat dataset.
@@
Table 6: Example prompt inputs from the web_questions dataset.
@@
We select $5$ widely-used LLMs as the specified models in our experiment, including Llama-2-13B [44], GPT-NeoX-20B [2], OPT-30B [52], Falcon-40B [1], and Llama-3.1-70B [9]. As alternative models, we use $6$ smaller LLMs, including GPT2-XL ($1.5$B) [36], GPT-NEO-2.7B [12], GPT-J-6B [47], OPT-6.7B [51], Vicuna-7B [56] and Llama-2-7B [44]. In Table 7, we list the number of parameters, hidden state dimension, and model developer for each LLM involved.
@@
F.2 Additional Protocol Training Details
@@
where the first item is the contrastive loss introduced in Eq. (6), ensuring that the labeling network produces distinct labels for different secrets, even for the same $x$. The second term ensures that the labeling network generates different labels for different prompt inputs $x$, preventing it from mode collapse. Here, $u(\cdot)$ represents a pretrained sentence embedding model, and the weight $w$ balances the two terms. We use all-mpnet-base-v2 [37] as the sentence embedding model and a 2-layer MLP to embed the secret. Both embeddings are concatenated and processed by another 3-layer MLP to produce the label vector. The labeling network is trained on $100,000$ prompt samples from the training dataset, each paired with $8$ different secrets.
@@
F.3 Experimental Details and Additional Results of the Protocol Accuracy
@@
If the hidden dimension of the alternative LLM, $d_{\mathcal{M}_{alt}}$, differs from that of the specified model, $d_{\mathcal{M}_{spec}}$, we apply a random projection matrix $W\in\mathbb{R}^{d_{\mathcal{M}_{alt}}\times d_{\mathcal{M}_{spec}}}$ to align the dimensions, where each element of $W$ is sampled from a standard normal distribution.
Cross-Dataset Generalization from a Fixed Training Set
To evaluate the generalizability of our protocol, we train the proxy task model and decision threshold solely on the LMSYS-Chat-1M dataset and assess performance on two unseen datasets. Specifically, we assess performance on the ToxicChat dataset [29], which contains toxic user prompts, and the web_questions dataset [24], which includes popular questions from real users. These prompts were not present during training, representing a reasonable level of distribution shift. As shown in Table 9 and 10, the FNR increases slightly for some models but remains within an acceptable range, while the FPR stays consistently low across various combinations of specified and alternative models. Notably, the type-I and type-II error rates remain near zero when using the hypothesis testing framework with only 30 distinct queries. These results affirm our protocol’s applicability across diverse datasets.
Training on Diverse Datasets
We further evaluate the versatility of our protocol by training and testing SVIP on two additional datasets: GSM8K [5] for mathematical reasoning and Verifiable-Coding-Problems [35] for program verification. Table 11 shows that SVIP maintains FPR below $3\%$ and FNR below $2\%$ on held-out test sets, demonstrating its strong performance across distinct task domains and confirming its applicability beyond conversational settings.
Table 9: FNR and FPR across different specified models on the ToxicChat dataset. The protocol is trained on the LMSYS-Chat-1M dataset.
Specified Model FNR $\downarrow$ FPR $\downarrow$
Llama-2-13B 3.40% 4.33% 3.65% 3.24% 4.21% 4.53% 5.12% 4.50%
GPT-NeoX-20B 15.35% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00%
OPT-30B 2.56% 0.00% 0.08% 0.12% 0.06% 0.18% 0.02% 0.04%
Falcon-40B 10.30% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00%
Llama-3.1-70B 9.24% 4.40% 5.83% 5.51% 6.12% 6.47% 5.27% 5.36%
Table 10: FNR and FPR across different specified models on the web_questions dataset. The protocol is trained on the LMSYS-Chat-1M dataset.
@@
Table 11: FNR and FPR across different specified models on two additional training datasets. The metrics are evaluated on held-out test sets.
Dataset Specified Model FNR $\downarrow$ FPR $\downarrow$
OPT-6.7B Llama-2-7B
GSM8K OPT-30B 2.28% 0.00% 0.00%
Llama-3.1-70B 1.03% 0.00% 0.00%
Verifiable-Coding-Problems OPT-30B 1.29% 1.31% 0.50%
Llama-3.1-70B 1.97% 1.40% 0.77%
Figure 8: Empirical distribution of the $L_{2}$ distance between the predicted proxy task output $f_{\phi^{*}}(z(x))$ and the label vector $y_{\gamma^{*}}(x,s)$ on the test dataset of LMSYS-Chat-1M for $2$ additional specified models.
F.4 Additional Results of the Computational Cost Analysis
For Table 12(a) and 12(b), all measurements were recorded on a single NVIDIA L40S GPU. Our protocol introduces minimal overhead for both the user and the computing provider during the deployment stage. Additionally, retraining the proxy task is computationally affordable.
Table 12: Computational costs of SVIP.
(a) Deployment stage costs.
Specified Model Runtime (Per Prompt Query) GPU Memory Usage
User Computing Provider User Computing Provider
Llama-2-13B 0.0056 s 0.0017 s
GPT-NeoX-20B 0.0057 s 0.0017 s
OPT-30B 0.0057 s 0.0018 s 1428 MB 980 MB
Falcon-40B 0.0057 s 0.0018 s
Llama-3.1-70B 0.0057 s 0.0019 s
(b) Proxy task retraining costs.
Specified Model Proxy Task Retraining Time
Llama-2-13B 4492 s
GPT-NeoX-20B 4500 s
OPT-30B 4580 s
Falcon-40B 4596 s
Llama-3.1-70B 5125 s
F.5 Examining the Labeling Network
@@
F.6 Experimental Details of Adapter Attack
Specifically, the attack succeeds if: $\|f_{\phi^{*}}(g_{{\theta}^{*}}(t_{\psi^{*}}(s)\oplus a_{{\lambda}^{*}}(h_{\mathcal{M}_{alt}}(x)))-y_{\gamma^{*}}(x,s)\|_{2}\leq\eta$. We experiment with $30$ independently sampled secrets, and report the average ASR on the test dataset as a function of the number of prompt samples collected. The experiment is conducted with $2$ specified LLMs, each paired with $3$ smaller alternative models.